OpenStudy (anonymous):

I NEED HELP WILL MEDAL PLEASE Roni recently bought a car for $22,000. Her research shows the car will depreciate by an average of $1,500 per year. If x represents the number of years Roni owns the car, which of the following equations represents the value of the car after x years? y = 22,000x - 1,500 y = 1,500x - 22,000 y = 22,000 + 1,500x y = 22,000 - 1,500x

OpenStudy (tkhunny):

Which one STARTS at 22000 and decreases by 1500 each year? y = 22,000x - 1,500 -- For x = 0, y = -1500 -- No good. y = 1,500x - 22,000 -- For x = 0, y = -22000 -- No good. y = 22,000 + 1,500x -- For x = 0, y = 22000 -- Maybe. y = 22,000 - 1,500x -- For x = 0, y = 22000 -- Maybe. Okay, considering the starting value, we've ruled out two bas results. Use the other information to rule out the last one.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

So the last one is correct?

OpenStudy (tkhunny):

You tell me. Does it meet ALL the criteria?

OpenStudy (anonymous):

i think so because if it is depreciating wouldnt it be subtracting?

OpenStudy (tkhunny):

Well, there you go. No more doubting. Move on to the next one.

OpenStudy (anonymous):

ok well my practice session just closed on me so the question is now irreverent but thanks for explaining to me better on how to do it